### v2

# High-resolution direct phase control in the spectral domain in ultrashort pulse lasers for pulse-shaping applications

Abstract
Ultrafast laser systems, those with a pulse duration on the order of picoseconds or less, have enabled advancements in a wide variety of fields. Of particular interest to this work, these laser systems are the key component to many High Energy Density (HED) physics experiments. Despite this, previous studies on the shape of the laser pulse within the HED community have focused primarily on pulse duration due to the relationship between pulse duration and peak intensity, while leaving the femtosecond scale structure of the pulse shape largely unstudied. To broaden the variety of potential pulses available for study, a method of reliably adjusting the pulse shape at the femtosecond scale using sub-nanometer resolution Direct Phase Control has been developed. This paper examines the capabilities of this new method compared to more commonplace dispersion-based pulse shaping methods. It also will detail the capabilities of the core algorithm driving this technique when used in conjunction with the WIZZLER and DAZZLER instruments that are common in high intensity laser labs. Finally, some discussion is given to possible applications on how the Direct Phase Control pulse shaping technique will be implemented in the future.
I Introduction
Ultrafast lasers, meaning laser systems with full width at half maximum (FWHM) durations on the order of ps or less, were first shown to be physically possible in the 1960s Bass and Woodward (1968); Cubeddu et al. (1969) However, the advent of Chirped Pulse Amplification in 1985 Strickland and Mourou (1985) was the catalyst that truly pushed the field forward, allowing for high intensity lasers with FWHM durations measured in fs to be made. Today, these fs class lasers can be found in labs around the world, with peak powers ranging from only a few kW all the way into the PW range. Pennington et al. (1997); Jeong and Lee (2014)
This wide range of power and increased availability has lead ultrafast lasers to be used in a wide variety of fields: from medical (LASIKShen et al. (2019) and optogeneticsJi et al. (2019)), to material processing Sugioka and Cheng (2014) and metrologyOh and Kim (2005); Udem, Holzwarth, and Hänsch (2002), to more pure research focused endeavors such as High Energy Density (HED) physicsDitmire et al. (2004). While each application brings with it a unique set of needs and challenges, this paper will focus mainly on applications for HED physics and the sensitivity these experiments have to the instantaneous field and intensity of the laser pulse, though the techniques and methods discussed may be useful to other fields and applications as well.
The difficulty in maintaining ultrashort pulses has increased as the pulse durations have shrank. This lead to the development of tools and systems to help control and stabilize the laser pulse. Critically to the work done in this paper, two instruments in particular have become fairly commonplace in ultrafast laser labs. The WIZZLER and the DAZZLER, both built by the company Fastlite, work hand in hand to maintain and adjust laser pulsesVerluise et al. (2000). Using a technique called Self-Referential Spectral Interferometry (SRSI), the WIZZLER is able to measure both the pulse spectral intensity and phase from a single shotOksenhendler et al. (2010), allowing it to produce a complete reconstruction of the shape of the pulse in time. Meanwhile, the DAZZLER is a type of Acousto-Optic Programmable Dispersive Filter (AOPDF)Tournois (1997) capable of selectively adjusting the phase of specific wavelengths relative to each other. When used together, the two instruments allow for a feedback loop wherein the WIZZLER measure the spectrum and phase before telling the DAZZLER how to adjust the phase to minimize the pulse duration.
The DAZZLER makes use of an acousto-optic, birefringent crystal to control relative spectral phase. Birefringent crystals have the unique property that the index of refraction along one axis (referred to as the ordinary axis) is different than the index of refraction along another axis (the extraordinary axis). Combined with the property of acousto-optic crystals to diffract light based on the interaction between the sound and electromagnetic waves passing through them, it is possible to have two different wavelengths of light have two different optical path lengths despite going through the same physical length of crystal. This difference in optical path lengths creates a difference in phase between the two wavelengths relative to each other. Thus, with a carefully crafted acoustic pulse going through the crystal, it is possible to add or remove spectral phase differences across the full bandwidth of a laser pulse in a single pass.
Refer to caption            LABEL:sub@fig:ftlRefer to caption
Refer to caption            LABEL:sub@fig:splitRefer to caption
Figure 1: Examples of pulse envelope and field superimposed over a sample of frequencies that make up the pulses. The black line is the electric field of the pulse (with pulse envelope plotted in gray around it), while the multi-colored plots are four different frequencies out of the many that were used to create the overlaid pulse. The amplitude of all frequencies were normalized to themselves for the purposes of demonstration. Both figures use the same spectrum, with LABEL:sub@fig:ftl is at FTL while LABEL:sub@fig:split has the phase of alternating frequencies shifted such that their peaks occur either slightly earlier or later than t=0.
The DAZZLER software calculates an acoustic pulse based on the requested phase input by the user. There are two ways the user can instruct the DAZZLER on how to adjust the phase. The first is by providing the DAZZLER with values for a 4th order dispersion polynomial, either through a settings file or the GUI interface for the DAZZLER software. This method has the benefit of being easily accessible and well understood, at the expense of being limited to phase profiles that can be described by a 4th order polynomial. The other option is Direct Phase Control (DPC), where an array of wavelengths, and the desired phase for each wavelength, are specified through a settings file. This is the same method that the WIZZLER uses during the feedback loop described earlier. DPC is more technically challenging as it changes the number of parameters the user is controlling from four to potentially several hundred, with each of these parameters having a different impact on the time varying amplitude of the pulse (commonly referred to as the pulse shape). However, DPC is also more powerful in terms of the variety of pulse shapes that it can create.
As mentioned earlier, the feedback loop between the WIZZLER and DAZZLER will attempt to get the pulse FWHM duration to be as close as possible to the minimum duration the system is capable of producing, a state referred to as Fourier Transform Limited (FTL). This is the default because FTL is generally the desired operating state for many current applications using ultrafast lasers. Intuitively, having the shortest duration possible means compressing the same amount of energy per pulse into the shortest amount of time, giving the highest peak power. Within the field of HED physics high peak power has been sought after due to the increase in the coupling efficiency between the laser and particle beams. However, recent researchZiegler et al. (2021); Khanna et al. (2024); Permogorov et al. (2022); Tayyab et al. (2018); Souri, Amrollahi, and Sadighi-Bonabi (2017); Kumar, Gopal, and Gupta (2019) has indicated that using specially shaped pulses may have benefits over having an FTL pulse.
Notably, ZeiglerZiegler et al. (2021) used a DAZZLER to add varying amounts of 3rd order dispersion (TOD) to their pulse for a proton acceleration experiment. From this, they observed an increased total number of protons as well as increased maximum energy for the protons compared to their results with an FTL pulse. They looked to quantify the link between these values and the numerical amount of TOD but did not attempt to explain the role of the pulse shape itself on the outcomes they observed. Similarly, PermogorovPermogorov et al. (2022) and TayyabTayyab et al. (2018) both looked at the effect of chirp on protons accelerated from thin foils. Tayyab did discuss how the changed pulse shapes due to the chirp they introduced could have influenced their result but did not make the pulse shape the focus of their work, while Permogorov intentionally isolated and controlled for pulse shape, making sure their laser pulse was symmetric. While SouriSouri, Amrollahi, and Sadighi-Bonabi (2017) and KumarKumar, Gopal, and Gupta (2019) both did look at the effect of the pulse shape using particle-in-cell simulations, no previous work could be found intentionally looking into the effect of laser pulse shape on laser driven plasma experimentally, creating a gap in the literature and leaving a potentially limited understanding of the underlying physics.
The work discussed herein begins to fill this gap in the present body of research by describing a method for implementing DPC over all spectral components, while also providing a reconstruction of the non-FTL pulse shape. This method relies on the initial spectrum of the pulse and its spectral phase. These two parameters were chosen because together they form a complete picture of the pulse and both can be measured simultaneously by the WIZZLER, providing the means to determine the spectrum and phase of a near-FTL pulse. Given the completeness of the pulse reconstruction, such a reconstruction is able to easily look into nearly any numerical aspect of the laser pulse, including, but not limited to, the intensity as a function of time, the instantaneous electric or magnetic field of the pulse, the chirp, and any phase changes. By using these near-FTL measurements as an initial point, it is also possible for an algorithm to provide an on-shot reconstruction of the pulse shape as DPC alters the spectral phase away from FTL, creating the opportunity for a robust diagnostic tool.
First, in section II briefly goes over the mathematical basis for pulse shaping. Section III then goes deeper into how the pulse reconstruction works while also explaining some of the capabilities and limitations of the hardware used for pulse shaping and testing. Section IV presents experimental data showcasing the performance of DPC pulse shaping. Section V summarizes the work and highlights some planned future developments.
II Mathematical Background
Mathematically, ultrashort laser pulses are nothing more than the superposition principle of light at work. If a large number waves, each with a unique frequency $\omega$, are superimposed with no restrictions imposed on the phase $\phi$ the resultant signal is purely noise since the waves will constructively and destructively interfere randomly. This resultant signal can be expressed as a sum of all the waveforms:
$E(t)=\sum_{\omega}A_{\omega}\exp{(-i\omega t+\phi)}+c.c.$ (1)
This is an application of the well-known Fourier series, written here in exponential form, with $A_{\omega}$ being the amplitude of each different frequency wave. If the phases of all the frequencies are all set to zero at any given time (for simplicity, time $t=0$) the resultant waveform will have one singular dominant peak with effectively zero signal outside this peak. This is the case for FTL laser pulses (figure 1a, pulse envelope showing one peak drawn in gray).
This dependence on phase is what allows the shape of the pulse to be controlled. By selectively shifting frequencies slightly away from zero phase, it is possible to change the overall shape of the pulse. Figure 1b shows exactly such a case. Each frequency has had the peak shifted in time by either plus or minus 0.58 fs from the center, alternating between plus and minus with each successive frequency. This is a non-physical system for the purposes of demonstrations. The spectrum being used is incredibly broad (600 nm, centered at 800 nm) and perfectly Gaussian across that bandwidth.
As can be seen in figure 1b, even this small shift of the peaks away from the center creates a symmetric double peaked pulse with minimal signal noise outside of the main pulse envelope. This type of pulse is very difficult to recreate using dispersion coefficients, as they struggle to create symmetric pulses with more than a single peak.
III Pulse Reconstruction
In order to numerically reconstruct the electric field of the laser pulse, the straightforward nature of the Fourier series is utilized. Using the spectral intensity and the spectral phase of the laser pulse, as measured by the WIZZLER, it is possible to perform a Discrete Fourier Transform (DFT) on the electric fields of the spectrum. This allows the spectral intensity and phase to be retrieved at any time while still providing a full description of the pulse. This method also allows for the flexible implementation of diagnostic features, such as the ability to input changes to the phase and have pulse parameters be quickly recalculated.
This process for pulse reconstruction is highly accurate and trustworthy, being identical to well-established theory. On top of that, results of the reconstruction algorithm developed for this method were compared to intensity measurements that the WIZZLER took for a near-FTL pulse and found the results matched to within rounding error (0.1 fs or less difference in FWHM duration). Qualitatively, the measured and reconstructed pulse shapes are clearly equivalent (figure 2). The data for this figure was taken using the GALADRIEL laser systemCollins et al. (2024) located at General Atomics in San Diego. This is a Ti:Sapphire laser with a center wavelength of 800 nm and a FWHM bandwidth of up to 80 nm.
Figure 2: A direct comparison of the pulse shape measured by the WIZZLER and reconstructed by the pulse reconstruction algorithm.
While this method of pulse shaping and reconstruction is independent of any particular system, it was designed with the intent of being used in partnership with a DAZZLER. To that end, it is worth briefly discussing the performance, and limitations, of the DAZZLER. The specific parameters of concern are the smallest spectral division the DAZZLER can effect, how much the DAZZLER can change the phase of each spectral step relative to the previous one, and how often can changes to the phase be sent to the DAZZLER such that those changes reliably affect the next laser pulse.
The first of these, what is essentially the spectral resolution of the DAZZLER, can be calculated based on system parameters. This relationship is given in the DAZZLER documentationdaz (2019) as:
$\delta\lambda_{1/2}=\frac{0.8\lambda^{2}}{\delta nL}\approx 8.9\frac{\lambda^{2}}{L}$ (2)
Here, $L$ is the length of the crystal in the DAZZLER system, $\delta n$ is the difference in index of refraction between the ordinary and extraordinary axes, and $\lambda$ is the center wavelength of the spectrum. The spectral resolution can also be found by looking at the settings files for the DAZZLER. For GALADRIEL, this value is 0.153 nm.
Refer to caption         LABEL:sub@fig:spolyphaseRefer to caption
Refer to caption         LABEL:sub@fig:spolyshapeRefer to caption
Refer to caption         LABEL:sub@fig:sdirephaseRefer to caption
Refer to caption         LABEL:sub@fig:sdireshapeRefer to caption
Figure 3: Examples of the capabilities of pulse shaping using dispersion polynomials compared to direct phase control. LABEL:sub@fig:spolyphase shows the spectrum and phase created using a dispersion polynomial while LABEL:sub@fig:spolyshape is the resultant pulse. Similarly, LABEL:sub@fig:sdirephase shows the spectrum and phase created using direct phase control with LABEL:sub@fig:sdireshape being the resultant pulse. Note that the spectrum used for both LABEL:sub@fig:spolyphase and LABEL:sub@fig:sdirephase is the same, only the phase is different.
The second parameter, the maximum phase change per spectral step, had to be found experimentally. The DAZZLER has limits built into its software that will prevent it from trying to output acoustic waveforms that it is unable to create. Based on those, the minimum number of steps needed to change the phase by $\pi$ radian was found to be 0.61 nm. In other words, when 800 nm was set to be zero phase, the next closest wavelength that could have a phase of $\pi$ radian was 800.61 nm, with the steps in between them set as linear steps between zero and $\pi$. This indicated that the maximum shift per division is roughly $\pi/4$ radian. It is possible this value is also system specific.
Similarly, testing on the GALADRIEL system has shown that the fastest frequency that the DAZZLER reliably change setpoints to have them accurately reflected in the next pulse is at 2 Hz. Trying to change setpoints more often than that would result in pulses occasionally not updating immediately, which gives incorrect data regarding pulse shape. The occurrence of this updating delay was not consistent, but would happen more often with more rapid attempts to send changes to the DAZZLER. In other words, it occurred on the order of once every hundred shots at 3 Hz, but on the order of once every 20 shots at 4 Hz, with no clear pattern that could be used to predict the next update delay.
Working within these parameters, demonstrating the difference in pulse shape control when using DPC as compared to using dispersion polynomials is straightforward. Using the same synthetic dataset used in figure 1 as a baseline, the phase was manipulated either with dispersion polynomials (fig. 3a) or direct phase control (fig. 3c). The goal was to transform the initial FTL pulse into pulse with two peaks that is as symmetric as possible. For the purposes of this work, these adjustments were done manually, while a future implementation of this technique will leverage machine learning to assist in building pulse shapes tailored at the fs level. For the dispersion polynomial control, 2nd through 4th order dispersion coefficients were used as those are what affect the pulse shape and are available in the DAZZLER GUI. To maintain parity with what can actually be done experimentally, the phase for the Direct Phase Control tests was not shifted by more than $\pi/4$ rad/step. Figures 3b and 3d show the pulse shapes that resulted from the respective phase changes.
IV Experimental Verification
In order for the Direct Phase Control pulse reconstruction to act as an on-shot diagnostic, a way for the pulse shape to be accurately reconstructed when other diagnostics, such as the WIZZLER, are unable to was needed. While baseline data can be easily recorded, certain experimental setups do not allow for extra diagnostics to be added close to the target. To enable these reconstructions, the core algorithm was designed to be flexible enough to recalculate the pulse reconstruction when provided new inputs such as a new spectral intensity measurements, changes to the provided spectral phase array, or specified dispersion coefficients. This flexibility enables the user to adjust the resultant pulse based on optical medium they know the laser will have to travel through, such as windows or dielectric coatings and still have an accurate reconstruction.
Refer to caption         LABEL:sub@fig:2ndorderRefer to caption
Refer to caption         LABEL:sub@fig:3rdorderRefer to caption
Refer to caption         LABEL:sub@fig:4thorderRefer to caption
Figure 4: A comparison of the predicted pulse shapes (in black) to ten measured pulses (colored). Each case started with the same FTL pulse and spectrum. Dispersion was then added to the spectral phase either computationally (for the DPC predicted pulse) or using the DAZZLER software GUI. The amount of dispersion added was as follows: LABEL:sub@fig:2ndorder 240 fs2 of 2nd order dispersion, LABEL:sub@fig:3rdorder 7500 fs3 of 3rd order dispersion, and LABEL:sub@fig:4thorder 240,000 fs4 of 4th order dispersion.
To show that the Direct Phase Control dispersion calculations accurately reflect changes to the laser pulse, a direct comparison with measured data was needed. This can be done by starting with an initial laser pulse optimized to be as close to FTL as possible using the standard WIZZLER-DAZZLER feedback loop. The DAZZLER GUI can then be used to introduce specific, known amounts of 2nd, 3rd, and 4th order dispersion to the pulse, while the WIZZLER is used to record both the initial and dispersion-added pulses. Feeding the same initial pulse and known dispersion values into the reconstruction algorithm will produce the expected pulse. For this testing, figure 4 shows the plots of the pulse intensity vs time for each of the dispersion orders. The black curve being the reconstructed pulse and the colored lines being a sampling of the data taken by the WIZZLER. Shot-to-shot measurment variation is illlurated by the different colored lines in Figure 4, and the reconstruction does a reasonable job within the measurement uncertainty. However, comparing the predicted pulse to each of these sampled pulses, the mean absolute error for the 2nd, 3rd, and 4th order dispersion was found to be $2.73\%$, $1.21\%$, and $2.38\%$ respectively.
Aiming to use the same method of comparison, the DAZZLER was provided with an array of specifically selected phases such that a dispersion polynomial wouldn’t be able to create (shown in the inset of figure 5). This was to show that the DAZZLER is capable of creating these arbitrary phase profiles and that it is possible to accurately predict the changes to pulse shape that result from DPC. The WIZZLER was used to record the initial and resultant pulses. The initial pulse spectrum and phase, as well as the specially adjusted phase array, were then used as inputs to the reconstruction algorithm. The results are plotted in figure 5. Again comparing the predicted pulse shape to each of the sampled pulses, the mean absolute error is $2.11\%$. It is worth noting that the WIZZLER is limited to measuring pulse shapes that are near-FTL (roughly seen as a FWHM that is 1.5 times FTL during testing, though this number was not thoroughly tested). However, the benchmarked algorithm used here has no such limitations and is capable of providing a pulse reconstruction on every shot regardless of shape or proximity to FTL.
Figure 5: A comparison of the predicted pulse shape (in black) compared to ten measured pulses (colored) after providing the DAZZLER with a custom phase array. The inset graph shows the spectrum (blue) and input spectral phase (red) for the laser pulse. Note the step function in the phase, which is difficult to replicate using a 4th order polynomial. The mismatch in the height of the two resultant pulses is believed to be due to the asymmetric laser spectrum on that day. With a perfectly square spectrum, the two resultant pulses should have been of roughly equal intensity.
V Conclusion and Next Steps
As it currently stands, the ability of this technique to accurately reconstruct precise pulse shape enables the creation of a reliable on-shot diagnostic tool. It is able to predict pulse shapes to within $<3\%$ error and will enable better on-shot information about the laser, and thus better inform the physics of what is happening in the laser plasma interaction. The implementation of Direct Phase Control pulse shaping on GALADRIEL is still being developed, with initial efforts on collecting data showing that the pulse becomes indistinct noise more rapidly than first expected. This is exacerbated by the WIZZLER’s inability to measure pulses that are too far from FTL. Future efforts for the study and implementation of this technique are planned using instruments that are more tolerant of non-FTL pulses but whose data analysis is likely too slow to be run for every shot at the  1-10 Hz repetition rate. Despite these limitations, DPC pulse shaping is capable of finely adjusting the pulse shape away from FTL and will unlock the bespoke laser pulses to probe yet unexplored phenomena.
Efforts in the near future will also see a form of this method integrated into the suite of diagnostics on the GALADRIEL system at General Atomics. It will be used for both on-shot diagnostics of the laser pulse as well as part of a control loop for the laser, helping improve stability for tests that involve pulse shaping. As a part of this control loop, it will eventually enable system users to specify arbitrary pulse shapes without needing to know the exact spectrum and phase needed to create them.
These capabilities rely on the assumption that the laser system is stable over some known amount of time. All lasers systems are known to have spectral drift over the course of a day of operation, but initial testing has shown that the GALADRIEL system remains close to an initial point over timescales on the order of single hours. Given that retaking baseline data and recalibrating takes only a couple of minutes, with possible upgrades to the lab planned to reduce that time even further, the amount of uptime is sufficient to justify the use of this method as a diagnostic given the significant advantages it provides.
VI Acknowledgements
S.B., M.M., and A.A. were supported in part by the National Science Foundation–Czech Science Foundation partnership (NSF Grant No. PHY-2206777).